### v2

# Bound States at Semiconductor–Mott Insulator Interfaces

Abstract
Utilizing the hierarchy of correlations in the context of a Fermi-Hubbard model, we deduce the presence of quasi-particle bound states at the interface between a Mott insulator and a semiconductor, as well as within a semiconductor–Mott–semiconductor heterostructure forming a quantum well. In the case of the solitary interface, the existence of bound states necessitates the presence of an additional perturbation with a minimal strength depending on the spin background of the Mott insulator. Conversely, within the quantum well, this additional perturbation is still required to have bound states while standing-wave solutions even exist in its absence.
I Introduction
The continuous advancements in thin film growth techniques of oxides [1, 2] have sparked a surge of interest in heterointerfaces between transition metal oxides [3, 4, 5, 6]. Electrons within these complex oxides experience a strong on-site Coulomb repulsion [7], giving rise to intricate electronic correlations while simpler oxides are band insulators or semiconductors. Interestingly, conducting interfaces between band-insulating oxide perovskites are well studied and understood, for example at the $\mathrm{LaAlO}_{3}/\mathrm{SrTiO}_{3}$ interface [8, 9].
The origin of the conducting layer is closely related to the avoidance of the so-called ’polar catastrophe’, a mechanism broadly applicable to (001) interfaces [10, 11, 12, 13, 14]. The polar heterojunction leads to charge transfer and subsequently to the formation and filling of a conduction band by the Ti $3d$ electrons [15]. Notably also the interface between the band insulator $\mathrm{KTaO}_{3}$ and Mott insulator $\mathrm{LaTiO}_{3}$ shows a conducting layer [16, 17]. In case of the (110) interface this cannot be traced back the ’polar catastrophe’, instead the Mott insulating physics becomes relevant [16].
Describing the strongly correlated nature of such Mott insulators necessitates specialized techniques such as strong-coupling perturbation theory [18, 19], dynamical mean-field theory [20] or methods like DFT+U [21], while semiconductors, on the other hand, can be accurately described by established band-structure methods, e.g. many-body perturbation theory.
The integration of these unlike materials into heterostructures holds the promise of unlocking novel applications [22, 23, 24, 25, 6]. Because of this, it is crucial to have a comprehensive understanding of the electronic states at their interfaces. While previous works have delved into the band lineup of heterointerfaces involving different Mott insulators and band insulators [26, 27] or calculated charge localization at specific Mott insulator interfaces like $\mathrm{LaAlO}_{3}/\mathrm{SrTiO}_{3}$ [28], $\mathrm{LaTiO}_{3}/\mathrm{KTaO}_{3}$ [29, 15] or $\mathrm{LaTiO}_{3}/\mathrm{SrTiO}_{3}$ [30, 31, 32], this study follows a general approach to deduce the existence of bound states in such type of systems, independent of specific band lineups and materials.
In navigating the complexities of these distinct material classes, we employ the hierarchy of correlations [33, 34, 33], a framework that aligns both weakly and strongly interacting materials on equal footing while preserving spatial resolution [35]. This approach allows us to bridge the gap between the different description requirements of semiconductors and Mott insulators, offering a unified perspective on their behavior within heterostructures. Within this approach, we will examine the different spin backgrounds of Mott insulators, the unpolarized background and the Mott-Néel state and their influence on charge bound states.
At first, we introduce the Hubbard Hamiltonian used to model both the Mott insulator and the semiconductor as well as the hierarchy of correlations. After this we calculate bound states at a single interface with an interface perturbation and give the minimal strength needed to support bound states for the different spin backgrounds. We continue with a heterostructure of an unpolarized Mott insulator stacked between two semiconductors with and without an interface perturbation.
II Hubbard Model and Hierarchy of Correlations
To characterize bound states at the interfaces of systems exhibiting varying degrees of correlation strength, we employ the Hubbard model [36], defined as follows:
$\hat{H}=-\frac{1}{Z}\sum_{\mu\nu s}T_{\mu\nu}c_{\mu s}^{\dagger}c_{\nu s}+\sum_{\mu}U_{\mu}\hat{n}_{\mu\uparrow}\hat{n}_{\mu\downarrow}+\sum_{\mu s}V_{\mu}\hat{n}_{\mu s}.$ (1)
Here, $c_{\mu s}^{\dagger}$ and $c_{\nu s}$ are the fermionic creation and annihilation operators, respectively, at sites $\mu$ and $\nu$. The corresponding number operators are denoted by $\hat{n}_{\mu s}$, where the spin index $s$ takes the values $\uparrow$ and $\downarrow$. The adjacency matrix $T_{\mu\nu}$ encodes both the lattice structure and hopping strength, in this work taking the form $T$ for nearest neighbors and zero otherwise. The coordination number $Z$ counts the nearest neighbors.
The parameters $U_{\mu}$ and $V_{\mu}$ serve to differentiate between strongly correlated systems ($U\neq 0$, $V=0$) and weakly correlated systems ($U=0$, $V\neq 0$), as the on-site repulsion $U_{\mu}$ and on-site potential $V_{\mu}$ play crucial roles in characterizing the nature of electronic correlations and the relative band alignments within the system. Note that a possibly small $U$ in a weakly correlated semiconductor could be treated by a mean-field approximation and hence simply be absorbed in the magnitude of an effective $V_{\mu}$. This is similar to Fermi liquid theory [37, 38].
II.1 Hierarchy of Correlations
To approximate solutions for charge modes at the interface, we employ the hierarchy of correlations [39, 34, 33] tailored for systems with a large coordination number $Z\gg 1$. The reduced density matrices of two or more lattice sites are decomposed into correlated and on-site components. Specifically, for two lattice sites $\mu$ and $\nu$, the decomposition reads $\hat{\rho}_{\mu\nu}=\hat{\rho}_{\mu}\hat{\rho}_{\nu}+\hat{\rho}_{\mu\nu}^{\text{corr}}$. Based on the large coordination number assumption $Z\gg 1$, we may employ an expansion into powers of $1/Z$, where we find that higher-order correlators are successively suppressed. The two-point correlator scales as $\hat{\rho}_{\mu\nu}^{\mathrm{corr}}=\mathcal{O}(Z^{-1})$, while the three-point correlation is suppressed as $\hat{\rho}_{\mu\nu\lambda}^{\mathrm{corr}}=\mathcal{O}(Z^{-2})$, and so on. This yields an iterative scheme to solve for the full density operator $\hat{\rho}$. More details are presented in App. A.
Similar to the idea of Hubbard X [40, 41] or composite operators [42], we introduce quasi-particle operators as:
$\hat{c}_{\mu sI}=\hat{c}_{\mu s}\hat{n}_{\mu\bar{s}}^{I}=\left\{\begin{array}[]{ccc}\hat{c}_{\mu s}(1-\hat{n}_{\mu\bar{s}})&{\rm for}&I=0,\\ \hat{c}_{\mu s}\hat{n}_{\mu\bar{s}}&{\rm for}&I=1\end{array}\right.$ (4)
for doublons $I=1$ and holes $I=0$. From these, we define the correlation functions $\langle\hat{c}_{\mu sI}\hat{c}_{\nu sJ}\rangle^{\text{corr}}$ [35]. For the relevant part of the dynamics, this can further be simlified by a factorization [43, 39], which yields the doublon $p_{\mu}^{1}$ and holon $p_{\mu}^{0}$ amplitudes as:
$\langle\hat{c}_{\mu sI}\hat{c}_{\nu sJ}\rangle^{\text{corr}}=(p_{\mu}^{I})^{*}p_{\nu}^{J}\,.$ (5)
The two amplitudes can be grouped together using a spinor notation and governing equations for these (quasi) particles [35] can be derived. In a hyper-cubic lattice dependent on the parallel momentum $k^{\|}$ they read:
$\left(E-U_{\mu}^{I}-V_{\mu}\right)p_{\mu}^{I}+\langle\hat{n}_{\mu}^{I}\rangle^{0}\sum_{J}T_{\bf k}^{\|}p_{\mu}^{J}$ (6)
$=-T\frac{\langle\hat{n}_{\mu}^{I}\rangle^{0}}{Z}\sum_{J}\left(p_{\mu-1}^{J}+p_{\mu+1}^{J}\right).$
The expectation values $\langle\hat{n}_{\mu}^{I}\rangle^{0}$ are to be taken in the mean-field background $\hat{\rho}^{0}_{\mu}$ and we use $U_{\mu}^{I}=IU_{\mu}$. More details on the derivation are presented in App. B.
II.2 Unpolarized Mean-Field Background
In the limit of strong correlation $U\gg T$, the mean-field state within the strongly correlated Mott insulator ensures one particle per site, with additional virtual hopping processes suppressed by $T^{2}/U^{2}$ [7, 44]. This charge background has different manifestations as there are different spin orderings with the same charge configuration. At first, there is the unpolarized background:
[…]
This might be created by a finite temperature destroying any spin ordering but to low to excite charges above the Mott gap $U$ or by spin frustration in lattices.
On the other hand, in the weakly correlated semiconductor region, where the quasi-particles are the real electrons, the on-site density matrix takes different forms for the valence and conduction bands: $\hat{\rho}_{\mu}^{0}=\ket{\uparrow\downarrow}_{\mu}\!\bra{\uparrow\downarrow}$ | for the valence band and $\hat{\rho}_{\mu}^{0}=\ket{0}_{\mu}\!\bra{0}$ | for the conduction band. Within each individual region, quasi-particles can be described using the ansatz $p_{\mu}^{I}=\alpha_{I}e^{i\kappa\mu}+\beta_{I}e^{-i\kappa\mu}$ [35], with the proportionality $Ep_{\mu}^{0}=(E-U)p_{\mu}^{1}$ in the Mott insulator. In the semiconducting sites, either one or the other is zero. The corresponding wave numbers are given by:
$\cos\kappa_{\mathrm{semi}}$ $=\frac{Z}{2T}\left[V-E-T_{\bf k}^{\|}\right],$ (8)
$\cos\kappa_{\mathrm{Mott}}$ $=\frac{Z}{2T}\left[\frac{E(U-E)}{E-U/2}-T_{\bf k}^{\|}\right].$
These expressions provide a description of quasi-particle behavior within the semiconductor and Mott insulator regions, offering valuable insights into their wave-like properties in these correlated systems [35]. In the following, we introduce the abbreviations:
$r_{\pm}=e^{\pm i\kappa_{\mathrm{Mott}}}\,\quad s_{\pm}=e^{\pm i\kappa_{\mathrm{semi}}}$ (9)
to describe the Mott and semiconducting solutions, respectively.
II.3 Bi-Partite Mean-Field Background
At second, there is the antiferromagnetic Mott-Néel state with its checkerboard structure:
[…]
In this spin configuration, the unit cell has double the size compared to the unpolarized background. This leads to backfolding in the Brillouin zone, giving two solutions: an even and an odd one. Moreover, the bi-partite structure will imprint on the weakly interacting sites.
To get this, we introduce another index $A$ or $B$ for the respective sublattice. In the Mott-Néel state, the coupled equations for doublons and holons thus read:
$Ep_{\mu}^{0_{A}}$ $=-\left[T_{\mathbf{k}}^{\|}p_{\mu}^{1_{B}}+\frac{T}{Z}\left(p_{\mu+1}^{1_{B}}+p_{\mu-1}^{1_{B}}\right)\right]$ (11)
$\left(E-U\right)p_{\mu}^{1_{B}}$ $=-\left[T_{\mathbf{k}}^{\|}p_{\mu}^{0_{A}}+\frac{T}{Z}\left(p_{\mu+1}^{0_{A}}+p_{\mu-1}^{0_{A}}\right)\right].$
Again, there is a proportionality between the two quasi-particles $p_{\mu}^{1_{B}}=\beta p_{\mu}^{0_{A}}$ with $\beta=\pm\sqrt{\frac{E}{E-U}}$. With the ansatz $p_{\mu}^{1_{B}}=\mathcal{B}\kappa^{\mu}$ and $p_{\mu}^{0_{A}}=\mathcal{A}\kappa^{\mu}$, and we find the eigenmodes:
$\kappa_{1,2}=$ $-\frac{Z}{2T}\left(\sqrt{E(E-U)}+T_{\mathbf{k}}^{\|}\right)$ (12)
±$\pm\sqrt{\left[\frac{Z}{2T}\left(\sqrt{E(E-U)}+T_{\mathbf{k}}^{\|}\right)\right]^{2}-1}$
$\kappa_{3,4}=$ $+\frac{Z}{2T}\left(\sqrt{E(E-U)}-T_{\mathbf{k}}^{\|}\right)$
±$\pm\sqrt{\left[\frac{Z}{2T}\left(\sqrt{E(E-U)}-T_{\mathbf{k}}^{\|}\right)\right]^{2}-1}$
with $\kappa_{1}\kappa_{2}=\kappa_{3}\kappa_{4}=1$. The proportionality between them reads
$B_{i}=\frac{1}{U-E}\left[T_{\mathbf{k}}^{\|}+\frac{T}{Z}\left(\kappa_{i}+\kappa_{i}^{-1}\right)\right]A_{i},$ (13)
which again fixes $\beta$. $\kappa_{1}$ and $\kappa_{2}$ belong to ’even’ solutions, i.e. having the same sign on neighboring sites of both sublattices, whereas $\kappa_{3}$ and $\kappa_{4}$ belong to ’odd’ solutions defined by a sign change between sublattices.
Without loss of generality, we can extend the bi-partite structure to the semiconducting half-space, now setting $p_{\mu}^{0_{A}}=\pm\alpha p_{\mu}^{0_{B}}$. Together with the ansatz $\lambda^{\mu}$, we find the eigenmodes for $\alpha=+1$ as:
$\lambda_{\pm}=-\frac{Z}{2T}\left(E-V+T_{\mathbf{k}}^{\|}\right)\pm\sqrt{\left[\frac{Z}{2T}\left(E-V+T_{\mathbf{k}}^{\|}\right)\right]^{2}-1}.$ (14)
This is again an ’even’ solution. Similar for $\alpha=-1$, we find with $p_{\mu}^{0_{A}}=\rho^{\mu}$
$\rho_{\pm}=+\frac{Z}{2T}\left(E-V-T_{\mathbf{k}}^{\|}\right)\pm\sqrt{\left[\frac{Z}{2T}\left(E-V-T_{\mathbf{k}}^{\|}\right)\right]^{2}-1}.$ (15)
belonging to the ’odd’ solution.
III Unpolarized Mott-Semiconductor interface
Figure 1: Energy schematics of the bound states at the single unpolarized interface. The solid line gives the relevant semiconductor band $V-T_{\mathbf{k}}$ while the dashed line depicts the two Mott bands around $E\approx 0$ and $E\approx U$. The dotted line describes the defining equation for the bound states that are shown as the thick vertical lines outside the bands.
First, our objective is a single unpolarized Mott-semiconductor interface. We consider a hypercubic lattice where the two half-spaces correspond to a Mott insulator with site index $\mu<0$ and a semiconductor with $\mu\geq 0$. In addition to the infinitely extended continuum states [35], bound states, characterized by their spatial decay away from the interface, may occur. For the complex numbers $r_{a}$ and $s_{b}$ in our ansatz, this implies
$p_{\mu}^{0}=\begin{cases}A_{a}\left(r_{a}\right)^{\mu}&\mu<0\text{ with }\left|r_{a}\right|>1\text{ in Mott},\\ B_{b}\left(s_{b}\right)^{\mu}&\mu\geq 0\text{ with }\left|s_{b}\right|<1\text{in semi.},\end{cases}$ (16)
where the indices $a,b$ stand for the $+$ or -- sign.
While we keep the model parameters $T_{\mu\nu},U_{\mu}$ and $V_{\mu}$ constant inside their respective regions, our analysis shows that an additional on-site potential $\Delta V\neq 0$ at the semiconductor interface site $\mu=0$ is a pre-requisite for obtaining interface states. This interface perturbation arises due to local modifications in the epitaxial interfaces of different materials. It might be introduced by a structural relaxation or deformation [45, 46, 47], by the tilting of the perovskite octahedra as in $\mathrm{LaTiO}_{3}/\mathrm{KTaO}_{3}$ [15] or other local modifications [48, 49].
By considering the relation between particles $p_{\mu}^{1}$ and holes $p_{\mu}^{0}$ in the Mott and semiconductor regions, we can establish boundary conditions at the interface using Eq. 6. These relate the amplitudes with the interface perturbation. After some algebra (see App. C for details), the defining equation for bound states at the single unpolarized interface reads
$\frac{1}{r_{a}}-\frac{1}{s_{b}}=\frac{\Delta VZ}{T}.$ (17)
It is important to note that this condition lacks a solution in the absence of an interface perturbation (see App. C.1). The defining equation is exemplarily shown in Fig. 1. The bound state energy solutions are marked. It looks similar to the Newns-Anderson model describing chemisorption or localized magnetic states in metals [50, 51]. Fig. 2 shows one solution to this equation, As expected for a bound state, the probability distribution exhibits localization at the interface – on the lattice site with the interface perturbation – and displays distinct decay constants towards both the Mott insulator and the semiconductor regions.
Figure 2: Bound state $p_{\mu}^{0}$ and quasi particle probability distribution $|p_{\mu}^{0}|^{2}$ at the interface for the unpolarized background with $E=1.16U$. The parameters are $V=1.1U$, $T=0.4U$ and $\Delta V=0.2U$.
III.1 Minimal Interface Perturbation
From an experimental standpoint, the selection of materials for the Mott insulator and the semiconductor will determine the interface perturbation $\Delta V$. Consequently, it is crucial to comprehend the requisite strength of the interface perturbation to establish a bound state. Interestingly, such a state is discovered to exist irrespective of whether $\Delta V$ is attractive or repulsive, provided it surpasses a specific threshold.
In order to calculate this threshold, we need to specify the allowed energy of a bound states. This is determined by the ansatz Eq. 16, which in the case $\Delta V>0$, is comprised of $r_{a}=r_{+}$ and $s_{b}=s_{-}$. To satisfy the condition $|s_{b}|=|s_{-}|<1$ any valid solution requires an energy $E>B=2T/Z-T_{\mathbf{k}}^{\|}+V$, ensuring that it lies energetically above the semiconducting band. To fulfill $|r_{+}|>1$, the upper band edges of the lower $A_{L}$ and of the upper $A_{H}$ Hubbard band are relevant:
$A_{L}$ $=\frac{1}{2}\left[-\sqrt{\left(T^{H}_{\mathbf{k}}\right)^{2}+U^{2}}+T^{H}_{\mathbf{k}}+2U\right],$ (18)
$A_{H}$ $=\frac{1}{2}\left[\sqrt{\left(T^{H}_{\mathbf{k}}\right)^{2}+U^{2}}+T^{H}_{\mathbf{k}}+2U\right],$
where $T^{H}_{\mathbf{k}}=\frac{2T}{Z}-T_{\mathbf{k}}^{\|}$.
Depending on the model parameters $V$ and $U$, four scenarios of band alignment arise. Firstly, the semiconducting band edge might be energetically higher than the Mott bands $B>A_{H}$, or equivalently, $V>U$ for $k^{\|}=0$. In this case, the threshold for the existence of bound states is given by Eq. 17 at $E=B$:
$\Delta V_{\mathrm{min,B}}=\frac{T}{Z}\left(\frac{1}{r_{+}(E=B)}-\frac{1}{s_{-}(E=B)}\right).$ (19)
In the limit $V\gg U$ this approaches $\Delta V_{\mathrm{min,B}}\to\frac{T}{Z}$. In this case, the required interface perturbation is small, $\Delta V\ll U$.
Secondly, for a semiconducting band edge situated between the center of the Mott gap and the upper Hubbard band $U/2<B<A_{H}$, the ansatz implies $E>A_{H}$. Consequently, Eq. 17 at $E=A_{H}$ yields:
$\Delta V_{\mathrm{min,A_{H}}}=\frac{T}{Z}\left(\frac{1}{r_{+}(E=A_{H})}-\frac{1}{s_{-}(E=A_{H})}\right).$ (20)
For strong Coulomb interaction compared to the hopping, this simplifies to $\Delta V_{\mathrm{min,A_{H}}}=\frac{1}{2}T_{\mathbf{k}}^{\|}+U-V+\mathcal{O}(T^{2})$.
Thirdly, if the semiconducting band edge is between the lower Hubbard band and the middle of the Mott band gap $A_{L}<B<U/2$, any bound state has $B<E<U/2$. The threshold is found as $\Delta V_{\mathrm{min,B}}$.
Fourthly, the semiconducting band edge might be energetically lower than the Hubbard bands $B<A_{L}$), which requieres $A_{L}<E$ and $A_{L}<E<U/2$ for bound states. Thus, Eq. 17 at $E=A_{L}$ gives:
$\Delta V_{\mathrm{min,A_{L}}}=\frac{T}{Z}\left(\frac{1}{r_{+}(E=A_{L})}-\frac{1}{s_{-}(E=A_{L})}\right).$ (21)
To first order in the hopping, this reads $\Delta V_{\mathrm{min,A_{L}}}=\frac{1}{2}T_{\mathbf{k}}^{\|}-V+\mathcal{O}(T^{2})$. For the full expressions in terms of the parameters $V$, $U$, $T$, and $T_{\mathbf{k}}^{\|}$ see App. C.2.
The required strength of the interface perturbation $\Delta V_{\mathrm{min}}$ depends on the band alignment. It is depicted in Fig. 3 (sold line) as a function of this band alignment, parametrized by the on-site potential $V$ for given hopping $T$. The approximate dependence of $T$, $U$, and $V$ is given. It decreases linearly with decreasing distance $V$ from the lower Hubbard band; is nearly constant in the lower half of the Mott gap; it goes up at $V=U/2$ and decreases from there linearly; before being constant again.
Figure 3: Minimally needed interface perturbation $\Delta V$ in the positive (solid) and negative (dashed) case as a function of the on-site potential $V$. The hopping strength $T=0.4U$ is fixed. The insets give the approximate formula for the minimal interface perturbation for $\Delta V>0$. The red dotted line gives the point symmetric rotated and shifted line of the positive interface perturbation onto the negative one.
Similar considerations as above might also be done for a negative interface perturbation $\Delta V<0$, corresponding to $r_{a}=r_{-}$ and $s_{b}=s_{+}$. In this case, any valid solution requires an energy below the semiconducting band edge $E<B_{L}=-2T/Z-T_{\mathbf{k}}^{\|}+V$, where both lower band edges in the Mott region, $A_{LL}$ and $A_{LH}$, are relevant. With $T^{L}_{\mathbf{k}}=2T/Z+T_{\mathbf{k}}^{\|}$ they read
$A_{LL}$ $=-\frac{1}{2}\left[T_{\mathbf{k}}^{L}-U+\sqrt{(T_{\mathbf{k}}^{L})^{2}+U^{2}}\right],$ (22)
$A_{LH}$ $=-\frac{1}{2}\left[T_{\mathbf{k}}^{L}-U-\sqrt{(T_{\mathbf{k}}^{L})^{2}+U^{2}}\right].$
The minimum required interface perturbation can be calculated using the same approach as before, and is shown in Fig. 3 (dashed line). The solutions for attractive and repulsive $\Delta V$ seem to be related (at least approximately) via an inflection point inside the Hubbard gap. To illustrate this, the red dotted line in Fig. 3 is obtained by a shift plus an inflection from the solid line. It nearly matches the dashed one, with the differences coming from the different band edges that need to be considered in the two different cases.
III.2 Decay Constant
The decay on both sides, $e^{-\kappa|\mu|}$, is characterized by an inverse length scale
$\kappa_{\mathrm{Mott}}$ $=\mathrm{arcosh}\left(\absolutevalue{\frac{Z}{2T}\left[\frac{E(U-E)}{E-U/2}-T_{\bf k}^{\|}\right]}\right),$ (23)
$\kappa_{\mathrm{semi}}$ $=\mathrm{arcosh}\left(\absolutevalue{\frac{Z}{2T}\left[V-E-T_{\bf k}^{\|}\right]}\right).$
The decay constant for the semiconducting side is illustrated in Fig. 4. Notably for small $T$, there exists a range of interface perturbations without any solutions (e.g., for $T=0.2U$, from $\Delta V/U\approx 0.5-1$).
Figure 4: The decay constant in the unpolarized semiconducting half-space is presented for various on-site potentials $V$ and hopping strengths $T$ across the interface perturbation $\Delta V$. Black symbols correspond to $V=1.1U$, blue symbols to $V=1.5U$, and red symbols to $V=0.2U$. Different shapes represent distinct hopping strengths: circles for $T=0.2U$, triangles for $T=0.5U$, and squares for $T=0.8U$.
Figure 5: The decay constant in the unpolarized Mott half-space is shown for different on-site potentials $V$ and hopping strengths $T$ over the interface perturbation $\Delta V$. Black symbols correspond to $V=1.1U$, blue to $V=1.5U$ and red to $V=0.2U$. The different symbol shapes depict different hopping strengths, circles for $T=0.2$, triangles for $T=0.5U$ and squares for $T=0.8U$.
Fig. 5 illustrates the same information for the Mott side. Again, for small hopping parameters (circles in Fig. 5), there is a range of $\Delta V$ values where no bound-state solution exists. In contrast to the semiconducting side, the decay constant in the Mott insulator may display non-monotonic behavior: When the on-site potential $V$ is close to the lower Hubbard band, a discontinuity occurs where $\kappa_{\mathrm{Mott}}$ sharply rises to a large value before abruptly decreasing close to zero. Increasing the interface perturbation for such a semiconducting band causes the energy of the solution to transition from being between the Hubbard bands to residing above the upper band, resulting in the observed discontinuity.
The rate of decay, whether it occurs more rapidly in the Mott or the semiconducting side, depends on both the on-site potential and the hopping strength. Fig. 6 displays both decay constants for an on-site potential close to the lower Hubbard band and another above the upper one, both for a small hopping strength. In the case of the larger on-site potential, the decay constant in the Mott is greater than in the semiconductor, indicating faster decay in the Mott. However, for the smaller on-site potential near the lower Hubbard band, there is a reversal as the interface perturbation increases. For small interface perturbations, the decay is faster in the Mott, but as the interface perturbation reaches a certain magnitude, the decay becomes slower than in the semiconductor.
Figure 6: Decay constant for both sides with $V=0.2U$ (black) and $V=1.5U$ (blue dashed) with $T=0.2U$ in the unpolarized case. The semi lines are on top of each other. Note the gap in the black curves.
Figure 7: Decay constant for a negatice interface perturbation $\Delta V$ in the unpolarized case. Circles give the $\kappa_{\mathrm{Mott}}$ and triangles give $\kappa_{\mathrm{semi}}$. The different colors depict different on-site potentials $V$. Hopping strength $T=0.2U$.
In the case of a negative interface perturbation, essentially all the characteristics observed for positive interface perturbations reappear, see Fig. 7.
Experimental measurements at the $\mathrm{LaTiO}_{3}/\mathrm{SrTiO}_{3}$ system show a tunnelling strength of $T=0.3\mathrm{eV}$ with a Coulomb repulsion of $U=6-20T$, such that realistic values are in the $T=0.16-0.05U$ regime. This matches the here used value of $T=0.2U$. The bound states found here have dispersion in the parallel direction. It is likely that all these subbands are partially filled leading to a metallic state parallel to the interface as it was measured in [1, 15, 32].
IV Unpolarized Mott region between two semiconductors
Utilizing the same methodology, we can also compute bound states in a system with two interfaces, where an unpolarized Mott insulator is positioned between two semi-infinite semiconductors, similar to a quantum well. Notably, there is no pre-requisite for the semiconductors to be identical; they can have distinct on-site potentials $V$ and interface perturbations $\Delta V$. The Mott region now spans $-a\leq\mu\leq a$, leading to the following ansatz:
$p_{n}^{0}=\begin{cases}As_{1}^{\mu}&\mu<-a\\ Br_{+}^{\mu}+Cr_{-}^{\mu}&-a\leq\mu\leq a\\ Bs_{2}^{\mu}&a<\mu\end{cases}$ (24)
Here, $|s_{1}|>1$ and $|s_{2}|<1$, with no requirement on the absolute value in the Mott region. The four boundary conditions now yield two conditions relating the amplitudes:
11 11 $=\frac{2(E-U)}{2E-U}\frac{s_{1}^{a+1}}{A}\left(Br_{+}^{-a-1}+Cr_{+}^{a+1}\right),$ (25)
11 11 $=\frac{2(E-U)}{2E-U}\frac{s_{2}^{-a-1}}{D}\left(Br_{+}^{a+1}+Cr_{+}^{-a-1}\right).$
Additionally, there are two defining equations for the interface perturbations at the interfaces:
$\frac{\Delta V_{1}Z}{T}$ $=\left[\frac{2(E-U)}{2E-U}\frac{Br_{+}^{-a}+Cr_{+}^{a}}{A}-s_{1}^{-a}\right]s_{1}^{a+1},$ (26)
$\frac{\Delta V_{2}Z}{T}$ $=\left[\frac{2(E-U)}{2E-U}\frac{Br_{+}^{a}+Cr_{+}^{-a}}{D}-s_{2}^{a}\right]s_{2}^{-a-1}.$
In the symmetric case, where the left and right semiconductors are the same, $s_{2}=s_{\bar{1}}$ holds to satisfy the condition of absolute values (with the bar denoting the opposite index).
Now, there exist even ($+$) solutions (with $B=C$, $A=D$) and odd (--) solutions (with $B=-C$, $A=-D$). With this, the boundary conditions for the coefficients and the defining equation for bound states in this system are given by:
11 11 $=\frac{2(E-U)}{2E-U}\frac{B}{A}\left(r_{+}^{-a-1}\pm r_{+}^{a+1}\right)s_{1}^{a+1},$ (27)
$\frac{\Delta VZ}{T}$ $=\frac{r_{+}^{a}\pm r_{+}^{-a}}{r_{+}^{a+1}\pm r_{+}^{-a-1}}-s_{1}.$
Here, $+$ denotes the even parity case, while -- indicates the odd one. Choosing $s_{1}=s_{\pm}$ alters the allowed energy range of the solutions to meet the requirement $|s_{1}|>1$; $s_{+}$ yields allowed energies above the semiconducting band, while $s_{-}$ yields those below.
IV.1 Standing-wave solutions
Figure 8: Defining equation with $\Delta V=-0.6U$, $T=0.4U$ and $V=1.5U$ in the zero parallel momentum sector and $a=6$. The crossings of the solid and dashed line with the dotted one give the energies for the even and odd parity solutions, respectively.
Disregarding any interface perturbation, the energy of the solutions must fall into one of the two Mott bands (see Fig. 8). Defining the Mott region as $\mu\in[-a,a]$, i.e. with $12a+$ 1 Mott lattice sites, we obtain $22a+$ 2 even-parity solutions from the defining equation. These are evenly distributed into $a+1$ solutions in both the lower and upper Hubbard bands. For odd parity, there are $2a$ solutions, equally divided into $a$ solutions in the lower and upper Hubbard bands. All of them have quasi-particle density leaking from the Mott into the semiconductors.
The fulfillment of the absolute value condition depends solely on the choice of $s_{1}$ and the potential $V$. A change in the potential has a minimal impact on the solutions, slightly lowering the energy (analogous to a potential well), as well as on the leaking. For solutions in the lower Hubbard band, the leakage is nearly independent of the on-site potential.
The solutions manifest as standing waves with $|r_{\pm}|=1$ in the Mott region. In both the lower and upper Hubbard bands, the solutions have energies $E_{0},E_{1},...,E_{2a}$, where even energies $E_{0}$, $E_{2}$, etc., correspond to even parity solutions, and odd energies $E_{1}$, $E_{3}$, and so forth, correspond to odd parity ones. The number of zeros of $p_{\mu}^{0}$ increases with energy from 0 to $2a$.
Enumerating only the even energy solutions as $E^{\mathrm{even}}_{0},E^{\mathrm{even}}_{1},\ldots E^{\mathrm{even}}_{a}$, the corresponding quasi-particle probability distributions $|p_{\mu}^{0}(E_{n}^{\mathrm{even}})|^{2}$ have $2(\frac{a}{2}-|\frac{a}{2}-n|)$ nodes. For $a=6$, this results in the sequence $0-2-4-6-4-2-0$. Performing the same analysis for the odd energy solutions $E^{\mathrm{even}}_{0},E^{\mathrm{even}}_{1},\ldots E^{\mathrm{even}}_{a-1}$, this yields $12(\frac{a}{2}-|\frac{a}{2}-n|)+$ 1 nodes ($1-3-5-5-3-1$ for $a=6$). In the lower Hubbard band the quasi-particle probability distribution is symmetric between states with the same number of nodes. In the upper Hubbard band, energies denoted by $E_{Un}$ ,the symmetry of the probability distribution is broken. As the energy of the solution increases, more of the probability becomes localized at the interfaces; with this the amount of leaked density increases as well.
Lowering the on-site potential $V$ such that the left band edge of the semiconducting band is below the lower Hubbard band results in the absence of $s_{1}=s_{-}$ solutions. Instead, they are replaced by $s_{1}=s_{+}$ solutions. For intermediate potentials with bands completely within the Mott gap, the lower Hubbard solutions are given by $s_{1}=s_{-}$, while the upper band ones are given by $s_{1}=s_{+}$. The energy of the solutions, as well as their $p_{\mu}^{0}$, remains unchanged when changing $s_{1}$ from $s_{-}$ to $s_{+}$; thus, bound state solutions persist within this heterostructure for all $V$.
As in the single interface case, the bound states display a dispersion in the parallel direction leading to subbands. These subbands are partially filled and as there is quasi particle weight at the interface the bound states exhibit metallic behavior there [1, 32, 15].
Figure 9: Bound states in the lower and upper Hubbard band with $\Delta V=0$, $T=0.4U$ and $V=1.5U$ in the zero parallel momentum sector and $a=6$.
IV.2 With interface perturbation
Introducing an interface perturbation results in one (or two in special cases) additional bound state(s) with $|r_{\pm}|\neq 1$. The interface perturbation must exceed a minimum value for these additional bound states to exist. It is independent of $s_{1}$ and given by
$\Delta V_{\mathrm{min}}=\frac{T}{Z}\left(-1+\frac{\cos(a\kappa_{\mathrm{Mott}}(E_{SL}))}{\cos((a+1)\kappa_{\mathrm{Mott}}(E_{SL}))}\right)$ (28)
with $E_{SL}=-2T/Z\pm T_{\mathbf{k}}^{\|}+V$. These additional bound states are illustrated in Fig. 10.
We note that two (rather than one) extra states emerge directly below and above one of the Hubbard bands if $V$ and $\Delta V$ align favorably. The lower state originates from the lowest solution initially inside the band being pushed outward.
The increase in the magnitude of the interface perturbation enhances localization at the interface, resulting in an increased decay constant in the semiconductors. The standing wave solutions within the Mott bands exhibit only mild sensitivity to the interface perturbation.
Figure 10: Additional bound state solutions for different interface perturbations $\Delta V=-0.25U$ and $-0.35U$. The latter one has two bound states, as the lowest standing wave solutions is pushed out of the band. $T=0.4U$ and $V=1.5U$ in the zero parallel momentum sector and $a=6$.
IV.3 Non-Symmetric Case
In the case of asymmetry, where $V_{1}\neq V_{2}$ and $\Delta V_{1}\neq\Delta V_{2}$, three out of the four equations Eq. 25, 26 govern the amplitudes $A$, $B$, $C$, and $D$, with one serving as the normalization constant. The remaining equation becomes the determining equation for the energy of the bound state solutions. These solutions are also asymmetric. In the absence of a interface perturbation, the half-space with the lower on-site potential, $V_{i}$, accumulates slightly more quasi-particle probability, typically by a single-digit percentage.
V Single Interface at Mott-Néel Background
At the antiferromagnetic Mott-Néel background, the interface bound states generally comprise both the even and odd solution. Choosing $\mu\leq 0$ as the semiconducting and $\mu>0$ as the Mott insulating sites, the ansatz is:
$\psi_{\mu\leq 0}=\begin{bmatrix}p_{\mu}^{0A}\\ p_{\mu}^{1A}\\ p_{\mu}^{0B}\\ p_{\mu}^{1B}\\ \end{bmatrix}=\frac{1}{2}A\lambda_{+}^{n}\begin{bmatrix}1\\ 1\\ 1\\ 1\end{bmatrix}+\frac{1}{2}B\rho_{-}^{n}\begin{bmatrix}1\\ 1\\ -1\\ -1\end{bmatrix}$ (29)
and
$\psi_{\mu>0}=\frac{1}{\sqrt{2}}\kappa_{i}^{n}\begin{bmatrix}A_{i}\\ 0\\ 0\\ B_{i}\end{bmatrix}+\frac{1}{\sqrt{2}}\kappa_{j}^{n}\begin{bmatrix}A_{j}\\ 0\\ 0\\ B_{j}\end{bmatrix}.$ (30)
Inserting this ansatz into the boundary conditions Eq. 62 yields:
$-\frac{Z}{T}\Delta V\frac{1}{2}(A-B)+A\lambda_{+}+B\rho_{-}$ $=\frac{1}{\sqrt{2}}B_{i}\kappa_{i}+\frac{1}{\sqrt{2}}B_{j}\kappa_{j}$ (31)
$A-B=$ $\frac{1}{\sqrt{2}}B_{i}+\frac{1}{\sqrt{2}}B_{j}$
$-\frac{Z}{T}\Delta V\frac{1}{2}(A+B)+A\lambda_{+}-B\rho_{-}$ $=\frac{1}{\sqrt{2}}A_{i}\kappa_{i}+\frac{1}{\sqrt{2}}A_{j}\kappa_{j}$
$A+B=$ $\frac{1}{\sqrt{2}}A_{i}+\frac{1}{\sqrt{2}}A_{j}$
These equations relate the interface perturbation $\Delta V$ to the amplitudes. As in the unpolarized case, there is a minimally needed strength of the additional interface perturbation for bound state to exist, but, in contrast, this threshold is constant at $\Delta V_{\mathrm{min}}\approx\frac{2T}{Z}$. Another distinct feature is that, for bound states with an energy below the semiconducting band, only a positive interface perturbation $\Delta V>0$ supports bound states. Even though the charge background with one electron per site is the same, the different spin backgrounds lead to distinctive behavior of the bound states.
VI Conclusions
Utilizing the hierarchy of correlations alongside the Fermi-Hubbard model we derived equations for the existence of bound states at single and multiple interface systems with different spin backgrounds. We account for an additional interface perturbation. These quasi-particle bound states manifest in the semiconductor as real electrons, while in the Mott insulator they are doublons and holons. Even though we used a model Hamiltonian, realistic values for the on-site potential, the hopping strength and the Coulomb repulsion can be extracted from literature.
At a solitary interface, the spin background and the charge backgrounds cannot be discussed independently: Both in the unpolarized case with (on average) one electron per site and in the antiferromagnetic Mott-Néel background an interface perturbation is needed to support bound states. The threshold of the perturbation strength depends on the spin background; in the unpolarized case it additionally depends on the band alignment, whereas in the Mott-Néel case it is constant.
In the case of a double-interface system, bound states exist even without an interface perturbation and manifest as standing wave solutions. The introduction of the interface perturbation adds additional states that are highly localized at the interface.
For future investigations, exploring second-order effects would be valuable. In the next hierarchical order of correlations, spin-spin and doublon-holon correlations emerge. They effectively renormalize the quasi particle energies used here, but they also potentially might have their own bound state structure. Additionally, studying the back reaction of correlation functions on the mean-field background could unveil insights, including potential space charge layer formation and effects of electron-density variation by spreading across the interface.
VII Acknowledgment
The authors thank F. Queisser and R. Schützhold for fruitful discussions and valuable feedback on the manuscript. This work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 278162697– SFB 1242.
Appendix A Hierarchy of Correlations
In order to describe the (quasi-)particles in the heterostructure, we resort to the hierarchy of correlations [39, 34, 33]. For any lattice Hamiltonian of the form:
$\hat{H}=\frac{1}{Z}\sum_{\mu\nu}\hat{H}_{\mu\nu}+\sum_{\mu}\hat{H}_{\mu}$ (32)
we can find an infinite set of equations for the density operator $\hat{\rho}$. In this, $\mu$ and $\nu$ are generalized coordinates. In the regime of large coordination number $Z\gg 1$ a truncation scheme is applicable to give a closed set of equations and an iterative way to solve this.
The starting point is the Heisenberg equation:
$i\partial_{t}\hat{\rho}=\big{[}H,\hat{\rho}\big{]}=\frac{1}{Z}\sum_{\mu\nu}\widehat{\mathcal{L}}_{\mu\nu}\hat{\rho}+\sum_{\mu}\widehat{\mathcal{L}}_{\mu}\hat{\rho}$ (33)
with the Liouville superoperators ^$\widehat{\mathcal{L}}_{\mu\nu}=\left[\hat{H}_{\mu\nu},\hat{\rho}\right]$ and ^$\widehat{\mathcal{L}}_{\mu}=\commutator{\hat{H}_{\mu}}{\hat{\rho}}$ ]. The next step is the decomposition of the density operator. Any operator whose expectation value one might be interested in is computed from a subset of lattice sites. This allows for the decomposition into on-site and correlated parts:
$\hat{\rho}_{\mu\nu}$ $=\hat{\rho}^{\text{corr}}_{\mu\nu}+\hat{\rho}_{\mu}\hat{\rho}_{\nu},$ (34)
$\hat{\rho}_{\mu\nu\lambda}$ $=\hat{\rho}^{\text{corr}}_{\mu\nu\lambda}+\hat{\rho}^{\text{corr}}_{\mu\nu}\hat{\rho}_{\lambda}+\hat{\rho}^{\text{corr}}_{\mu\lambda}\hat{\rho}_{\nu}+\hat{\rho}^{\text{corr}}_{\nu\lambda}\hat{\rho}_{\mu}+\hat{\rho}_{\mu}\hat{\rho}_{\nu}\hat{\rho}_{\lambda},$
and so on. In order to get the time evolution of these, we need to calculate:
$i\partial_{t}\hat{\rho}_{\mu}=\frac{1}{Z}\sum_{\alpha\neq\mu}\mathrm{tr}_{\alpha}\left(\widehat{\mathcal{L}}^{S}_{\alpha\mu}\left[\hat{\rho}_{\mu\alpha}^{\text{corr}}+\hat{\rho}_{\alpha}\hat{\rho}_{\mu}\right]\right)+\widehat{\mathcal{L}}_{\mu}\hat{\rho}_{\mu}.$ (35)
with the symmetrized from ^$\widehat{\mathcal{L}}^{S}_{\mu\nu}=\widehat{\mathcal{L}}_{\mu\nu}+\widehat{\mathcal{L}}_{\nu\mu}$ for the on-site density operator. This time evolution contains the two-point correlator $\hat{\rho}_{\mu\nu}^{\mathrm{corr}}$. The same is done for the two-site density operator, such that we can combine these two results to find:
$i\partial_{t}\hat{\rho}^{\text{corr}}_{\mu\nu}=$$\widehat{\mathcal{L}}_{\mu}\hat{\rho}^{\text{corr}}_{\mu\nu}+\frac{1}{Z}\widehat{\mathcal{L}}_{\mu\nu}\left(\hat{\rho}^{\text{corr}}_{\mu\nu}+\hat{\rho}_{\mu}\hat{\rho}_{\nu}\right)$ (36)
[…]
$+(\mu\leftrightarrow\nu)$
as the time evolution of the two-point correlator $\hat{\rho}_{\mu\nu}^{\mathrm{corr}}$. It contains the three-point correlator $\hat{\rho}_{\mu\nu\lambda}^{\mathrm{corr}}$. This builds up a set of equations:
$i\partial_{t}\hat{\rho}_{\mu}$ $=$ $F_{1}(\hat{\rho}_{\mu},\hat{\rho}_{\mu\nu}^{\mathrm{corr}})\,,$
$i\partial_{t}\hat{\rho}_{\mu\nu}^{\mathrm{corr}}$ $=$ $F_{2}(\hat{\rho}_{\mu},\hat{\rho}_{\mu\nu}^{\mathrm{corr}},\hat{\rho}_{\mu\nu\lambda}^{\mathrm{corr}})\,,$
$i\partial_{t}\hat{\rho}_{\mu\nu\lambda}^{\mathrm{corr}}$ $=$ $F_{3}(\hat{\rho}_{\mu},\hat{\rho}_{\mu\nu}^{\mathrm{corr}},\hat{\rho}_{\mu\nu\lambda}^{\mathrm{corr}},\hat{\rho}_{\mu\nu\lambda\kappa}^{\mathrm{corr}})\,,$
$i\partial_{t}\hat{\rho}_{\mu\nu\lambda\alpha}^{\mathrm{corr}}$ $=$ $F_{4}(\hat{\rho}_{\mu},\hat{\rho}_{\mu\nu}^{\mathrm{corr}},\hat{\rho}_{\mu\nu\lambda}^{\mathrm{corr}},\hat{\rho}_{\mu\nu\lambda\kappa}^{\mathrm{corr}},\hat{\rho}_{\mu\nu\lambda\kappa\beta}^{\mathrm{corr}}).$ (37)
The specific form of the functionals $F_{n}$ is dictated by the Hamiltonian.
If the initial state satisfies scaling relations such that $\ell$-point correlations are of order $\mathcal{O}(Z^{-\ell+1})$, this scaling persists for all times [52, 39]. Exploiting the scaling behavior, we approximate the equations to zeroth and first order, yielding:
$i\partial_{t}\hat{\rho}_{\mu}$ $\approx F_{1}(\hat{\rho}_{\mu},0),\quad\text{with solution}\quad\hat{\rho}_{\mu}^{0},$ (38)
$i\partial_{t}\hat{\rho}_{\mu\nu}^{\text{corr}}$ $\approx F_{2}(\hat{\rho}_{\mu}^{0},\hat{\rho}_{\mu\nu}^{\text{corr}},0).$
These two equations are used to describe the charge modes within the studied system. $\hat{\rho}_{\mu}^{0}$ encodes the mean-field background charge and spin structure.
Appendix B Hierarchy for the Fermi Hubbard Model
In order to apply the hierarchy of correlations to the Fermi-Hubbard model Eq. 1, we first introduce the two different spin backgrounds. There is the unpolarized state:
[…]
and there is the antiferromagnetic Mott-Néel state with its two sublattices $A$ and $B$ arranged in a checkerboard structure:
[…]
Independent of the mean-field background, it is instructive to introduce quasi-particle operators, similar to the idea of Hubbard X [40, 41] or composite operators [42], as:
$\hat{c}_{\mu sI}=\hat{c}_{\mu s}\hat{n}_{\mu\bar{s}}^{I}=\left\{\begin{array}[]{ccc}\hat{c}_{\mu s}(1-\hat{n}_{\mu\bar{s}})&{\rm for}&I=0,\\ \hat{c}_{\mu s}\hat{n}_{\mu\bar{s}}&{\rm for}&I=1\end{array}\right.$ (43)
for doublons $I=1$ and holes $I=0$. These better describe the physics, but note that these operators are approximately, but not exactly equal to the quasi-particle creation and annihilation operators for holons and doublons, see, e.g. [53]. The label $\bar{s}$ denoted the spin index opposite to $s$. For the correlation functions $\langle\hat{c}_{\mu sI}\hat{c}_{\nu sJ}\rangle$ we find:
$i\partial_{t}\langle\hat{c}^{\dagger}_{\mu sI}\hat{c}_{\nu sJ}\rangle^{\mathrm{corr}}=\frac{1}{Z}\sum_{\lambda L}T_{\mu\lambda}\langle\hat{n}_{\mu\bar{s}}^{I}\rangle^{0}\langle\hat{c}^{\dagger}_{\lambda sL}\hat{c}_{\nu sJ}\rangle^{\mathrm{corr}}$ (44)
$-\frac{1}{Z}\sum_{\lambda L}T_{\nu\lambda}\langle\hat{n}_{\nu\bar{s}}^{J}\rangle^{0}\langle\hat{c}^{\dagger}_{\mu sI}\hat{c}_{\lambda sL}\rangle^{\mathrm{corr}}$
$+\left(U_{\nu}^{J}-U_{\mu}^{I}+V_{\nu}-V_{\mu}\right)\langle\hat{c}^{\dagger}_{\mu sI}\hat{c}_{\nu sJ}\rangle^{\mathrm{corr}}$
$+\frac{T_{\mu\nu}}{Z}\left(\langle\hat{n}_{\mu\bar{s}}^{I}\rangle^{0}\langle\hat{n}_{\nu s}^{1}\hat{n}_{\nu\bar{s}}^{J}\rangle^{0}-\langle\hat{n}_{\nu\bar{s}}^{J}\rangle^{0}\langle\hat{n}_{\mu s}^{1}\hat{n}_{\mu\bar{s}}^{I}\rangle^{0}\right)+\mathcal{O}(1/Z^{2})$
Here, we used the abbreviation $U_{\mu}^{I}=IU_{\mu}$, i.e., $U_{\mu}^{I}=U_{\mu}$ for $I=1$ and $U_{\mu}^{I}=0$ for $I=0$.
The relevant part of this time evolution, which describes the dynamics, can further be simplified by a factorization [43, 39]. This yields the doublon ($I,J=1$) and holon ($I,J=0$) amplitudes:
$\langle\hat{c}_{\mu sI}\hat{c}_{\nu sJ}\rangle^{\text{corr}}=(p_{\mu}^{I})^{*}p_{\nu}^{J}\,.$ (45)
The two amplitudes can be grouped together using a spinor notation and governing equations for these (quasi) particles [35] can be derived. Assuming a highly symmetric lattice such as a hyper-cubic one allows us to perform the corresponding Fourier transform parallel to the interface by:
$p_{\mu sI}$ $=\frac{1}{\sqrt{N^{\parallel}}}\sum_{\mathbf{k}^{\parallel}}p_{n,\mathbf{k}^{\parallel},s}^{I}e^{i\mathbf{k}^{\parallel}\cdot\mathbf{x}_{\mu}^{\parallel}},$ (46)
$T_{\mu\nu}$ $=\frac{Z}{N^{\parallel}}\sum_{\mathbf{k}^{\parallel}}T_{m,n,\mathbf{k}^{\parallel}}e^{i\mathbf{k}^{\parallel}\cdot\left(\mathbf{x}_{\mu}^{\parallel}-\mathbf{x}_{\nu}^{\parallel}\right)}$
For the isotropic nearest neighbor hopping $T^{\|}_{n}=T^{\perp}_{n,n-1}=T$ the components read:
$T_{m,n,\mathbf{k}^{\parallel}}$ $=\frac{T^{\parallel}_{\mathbf{k}^{\parallel}}}{Z}\delta_{m,n}+\frac{T}{Z}(\delta_{n,n-1}+\delta_{n,n+1}),$ (47)
$T_{\mathbf{k}^{\parallel}}^{\parallel}$ $=2T\sum_{x_{i}}\cos(p_{x_{i}}^{\parallel})\equiv ZT_{\mathbf{k}}^{\|}\,,$
with the hopping contribution $T_{\bf k}^{\|}$. With this, the doublon and holon amplitudes follow the coupled equations:
$\left(E-U_{\mu}^{I}-V_{\mu}\right)p_{\mu}^{I}+\langle\hat{n}_{\mu}^{I}\rangle^{0}\sum_{J}T_{\bf k}^{\|}p_{\mu}^{J}$ (48)
$=-T\frac{\langle\hat{n}_{\mu}^{I}\rangle^{0}}{Z}\sum_{J}\left(p_{\mu-1}^{J}+p_{\mu+1}^{J}\right).$
The different possibilities of the mean-field background contribute by the expectation values $\langle\hat{n}_{\mu}^{I}\rangle^{0}=\text{Tr}(\hat{n}_{\mu}^{I}\hat{\rho}_{\mu}^{0})$. In the unpolarized state we have $\langle\hat{n}_{\mu}^{I}\rangle^{0}=1/2$, while in the antiferromagnetic Mott-Néel state we need another index for the $A$- and $B$-sublattice as one of these has $\langle\hat{n}_{\mu_{X}\uparrow}\rangle=0$ and $\langle\hat{n}_{\mu_{X}\downarrow}\rangle=1$, while the other one has them reversed $\langle\hat{n}_{\mu_{Y}\uparrow}\rangle=1$, $\langle\hat{n}_{\mu_{Y}\downarrow}\rangle=0$.
Appendix C Single Unpolarized Interface
At the single interface between the Mott insulator $\mu<0$ and the semiconductor $\mu\geq 0$ we can write down two boundary conditions from Eq. 48 with $\langle\hat{n}_{\mu<0}^{I}\rangle^{0}=1/2$, $\langle\hat{n}_{\mu\geq 0}^{1}\rangle^{0}=0$ and $\langle\hat{n}_{\mu\geq 0}^{0}\rangle^{0}=1$. Together with the relation between particles and holes $p_{\mu<0}^{0}E=p_{\mu<0}^{1}(E-U)$ in the Mott, the boundary conditions read:
$\left(E-V-\Delta V+T_{\mathbf{k}}^{\|}\right)p_{0}^{0}=$ $-\frac{T}{Z}\left(p_{1}^{0}+\frac{2E-U}{2(E-U)}p_{-1}^{0}\right),$ (49)
$\left(E+\frac{2E-U}{2(E-U)}T_{\mathbf{k}}^{\|}\right)p_{-1}^{0}=$ $-\frac{T}{Z}\left(p_{0}^{0}+\frac{2E-U}{2(E-U)}p_{-2}^{0}\right).$
Additionally, two identities hold within each individual region:
$E+T_{\mathbf{k}}^{\|}\frac{2E-U}{2(E-U)}$ $=-\frac{T}{Z}\left(r_{a}+\frac{1}{r_{a}}\right)\frac{2E-U}{2(E-U)},$ (50)
$\left(E-V+T_{\mathbf{k}}^{\|}\right)$ $=-\frac{T}{Z}\left(s_{b}+\frac{1}{s_{b}}\right).$
By combining the ansatz Eq. 16, the boundary conditions Eq. 49, and using the identities Eq. 50, we derive a relation between the amplitudes:
$\frac{2E-U}{2(E-U)}\frac{A_{a}}{B_{b}}=1.$ (51)
Furthermore, we obtain an equation relating the interface perturbation to the amplitudes and half-space solutions $r_{a}$ and $s_{b}$:
$\frac{\Delta VZ}{T}=\frac{2E-U}{2(E-U)}\frac{A_{a}}{B_{b}}\frac{1}{r_{a}}-\frac{1}{s_{b}}.$ (52)
These two equations combine to the defining equation for bound states at the unpolarized single interface Eq. 17.
C.1 Defining Equation
The defining equation from these two then reads:
$0=\frac{1}{r_{a}}-\frac{1}{s_{b}}-\frac{\Delta VZ}{T}.$ (53)
Without an interface perturbation, $\Delta V\equiv 0$, this equation is solved by $b\kappa_{\mathrm{semi}}=a\kappa_{\mathrm{Mott}}+2\pi c$ with an integer $c\in\mathbb{Z}$. But, this would also result in $|e^{ai\kappa_{\mathrm{Mott}}}|=|e^{bi\kappa_{\mathrm{semi}}}|$ which is in conflict with the ansatz Eq. 16. So, there are no solutions without the interface perturbation. Inserting $r_{a}$ and $s_{b}$ yields a function
$f(E)$ $=$ $\frac{1}{r_{a}}-\frac{1}{s_{b}}-\frac{\Delta VZ}{T}$ (54)
$=$ $ai\sqrt{1-\frac{Z^{2}\left(2E\left(\frac{E}{U-2E}+1\right)+T_{\mathbf{k}}^{\|}\right)^{2}}{4T^{2}}}+$
$+bi\sqrt{1-\frac{Z^{2}(E+T_{\mathbf{k}}^{\|}-V)^{2}}{4T^{2}}}$
$+\frac{Z(E(U-2V)+UV)}{4ET-2TU}-\frac{\Delta VZ}{T}$
whose zeros give the bound states. $a$ and $b$ are the $+$ or --, respectively. The imaginary “bells” coincide with the Mott and semiconductor band, respectively, and are not affected by the interface perturbation. The term $-\Delta VZ/T$ shifts the real part of the defining equation up and down. In Fig. 11 the defining equation for $r_{a}=r_{+}$ and $s_{b}=s_{-}$ is shown without and with a interface perturbation together with the absolute value $|r_{a}|$ and $|s_{b}|$. Without any interface perturbation the zeros of the real part are inside the non-zero imaginary part. So, by adding the interface perturbation it follows directly that any solution is outside of the bands and has imaginary $\kappa_{\mathrm{Mott}}$ and $\kappa_{\mathrm{semi}}$. The conditions $|r_{a}|>1$ and $|s_{b}|<1$ then dictate on which side of the “bells” the solution needs to be, because the absolute value (at least for the semiconductor) is a monotone function of energy $E$ interrupted by the plateau of the band. For the Mott, this is also true, but at $E=U/2$ there the absolute value goes down to zero again.
Figure 11: Eq. 54 with $r_{a}=r_{+}$ and $s_{b}=s_{-}$. The solid line gives the real part without any interface perturbation $R_{0}=\mathrm{Re}f(E,\Delta V=0)$, the dashed line the imaginary part $I=\mathrm{Im}f(E)$ and the dotted the real part with non-zero interface perturbation $R_{1}=\mathrm{Re}f(E,\Delta V=0.2U)$. The parameters are $V=1.1U$, $T=0.4U$
C.2 Minimum Interface Perturbation
The minimum required interface perturbation to shift the real part depends on the band alignment of the semiconducting band relative to the Hubbard bands. Mathematically, it depends on the on-site potential $V$ as this shifts the band edge. Now, one needs to distinguish four different cases: the semiconductor band edge is i) above of the upper Hubbard band; ii) between the band gap center $U/2$ and the upper Hubbard band; iii) above the lower Hubbard band but below the band gap center; iv) below the lower Hubbard band. In the first case, the minimum interface perturbation reads:
$4\Delta V_{\mathrm{min,B}}=\frac{U^{2}Z}{4T-Z(2T_{\mathbf{k}}^{\|}+U-2V)}+U-2V$
$-\frac{4iT\sqrt{1-\frac{\left(\alpha+\beta\right)^{2}}{4T^{2}(Z(2T_{\mathbf{k}}^{\|}+U-2V)-4T)^{2}}}}{Z}$ (55)
with the abbreviations:
$\alpha$ $=8T^{2}-4TZ(T_{\mathbf{k}}^{\|}+U-2V)$ (56)
$\beta$ $=Z^{2}\left(-2V(T_{\mathbf{k}}^{\|}+U)+T_{\mathbf{k}}^{\|}U+2V^{2}\right)$
In the second case, we find:
$4Z\Delta V_{\mathrm{min,A_{H}}}=\sqrt{4T^{2}-4TT_{\mathbf{k}}^{\|}Z+Z^{2}\left((T_{\mathbf{k}}^{\|})^{2}+U^{2}\right)}$
$-4iT\sqrt{1-\frac{\left(\alpha+2T+Z(T_{\mathbf{k}}^{\|}+U-2V)\right)^{2}}{16T^{2}}}$
$-2T+Z(T_{\mathbf{k}}^{\|}+U-2V)$ (57)
with the abbreviation:
$\alpha=\sqrt{(2T-ZT_{\mathbf{k}}^{\|})^{2}+Z^{2}U^{2}}$ (58)
as the minimum interface perturbation. In the third case, we get $\Delta V_{\mathrm{min,B}}$ again. In the fourth case we find:
$4Z\Delta V_{\mathrm{min,A_{H}}}=-\sqrt{4T^{2}-4TT_{\mathbf{k}}^{\|}Z+(T_{\mathbf{k}}^{\|})^{2}Z^{2}+U^{2}Z^{2}}$
$-4iT\sqrt{1-\frac{\left(-\alpha+2T+Z(T_{\mathbf{k}}^{\|}+U-2V)\right)^{2}}{16T^{2}}}$
$-2T+T_{\mathbf{k}}^{\|}Z+UZ-2VZ$ (59)
with:
$\alpha=\sqrt{(2T-ZT_{\mathbf{k}}^{\|})^{2}+Z^{2}U^{2}}$ (60)
C.3 Single Interface at Mott-Néel Background
Without loosing any generality, we use the mean-field state Eq. 10. This fixes the occupation number expectation values to $\langle\hat{n}_{\mu_{A}\uparrow}\rangle=1$, $\langle\hat{n}_{\mu_{B}\uparrow}\rangle=0$. Combining Eq. 48 with the ansatz Eq. 29 and Eq. 30 yields four boundary conditions:
$(E-V-\Delta V)p_{0}^{0A}$ $=-\left[T_{\mathbf{k}}^{\|}p_{0}^{0B}+\frac{T}{2Z}p_{1}^{1B}+\frac{T}{Z}p_{-1}^{0B}\right],$ (61)
$Ep_{1}^{0A}$ $=-\left[T_{\mathbf{k}}^{\|}p_{1}^{1B}+\frac{T}{Z}p_{2}^{1B}+\frac{2T}{Z}p_{0}^{0B}\right],$
$(E-V-\Delta V)p_{0}^{0B}$ $=-\left[T_{\mathbf{k}}^{\|}p_{0}^{0A}+\frac{T}{2Z}p_{1}^{0A}+\frac{T}{Z}p_{-1}^{0A}\right],$
$(E-U)p_{1}^{1B}$ $=-\left[T_{\mathbf{k}}^{\|}p_{1}^{0A}+\frac{T}{Z}p_{2}^{0A}+\frac{2T}{Z}p_{0}^{0A}\right].$
By using Eq. 11 within the single regions these might be simplified to:
$-\frac{Z}{T}\Delta Vp_{0}^{0AS}+2p_{1}^{0BS}$ $=p_{1}^{1BM},$ (62)
$2p_{0}^{0BS}$ $=p_{0}^{1BM},$
$-\frac{Z}{T}\Delta Vp_{0}^{0BS}+2p_{1}^{0AS}$ $=p_{1}^{0AM},$
$2p_{0}^{0AS}$ $=p_{0}^{0AM}.$
The additional index $M$ or $S$ marks the Mott and semiconducting spinor, respectively.